The Warren Watch Company sells watches for $21, fixed costs are $100,000, and variable costs are $11 per watch.

  1. What is the firm's gain or loss at sales of 6,000 watches? Enter loss (if any) as negative value. Round your answer to the nearest cent.
    $

    What is the firm's gain or loss at sales of 17,000 watches? Enter loss (if any) as negative value. Round your answer to the nearest cent.
    $

  2. What is the break-even point (unit sales)? Round your answer to the nearest whole number.
    units

  3. What would happen to the break-even point if the selling price was raised to $32?
    -Select-The result is that the break-even point is higher. The result is that the break-even point remains unchanged. The result is that the break-even point is lower. Item 4

  4. What would happen to the break-even point if the selling price was raised to $32 but variable costs rose to $23 a unit? Round your answer to the nearest whole number.


    -Select-The result is that the break-even point decreases. The result is that the break-even point remains unchanged. The result is that the break-even point increases. Item 5
Answer #1

1.
=6000*(21-11)-100000=-40000

2.
=17000*(21-11)-100000=70000

3.
=100000/(21-11)=10000

4.
=100000/(32-11)=4761.90476190476
The result is that the break-even point is lower

5.
=100000/(32-23)=11111.1111111111
The result is that the break-even point increases.